OpenStudy (anonymous):

Which fraction represents the probability of rolling a 1, 2, 4, or 5 when using a number cube with the numbers 1, 2, 3, 4, 5, and 6?

OpenStudy (lucaz):

the possible outcomes are 1, 2, 3, 4, 5, 6, total of 6, you want the probability of getting 1, 2, 4, 5, total of 4, P=4/6=2/3